OpenStudy (anonymous):

MEDAL AND A FAN i need help with 8 different question about writing linear functions. I will open up a new question for each so you can get 8 medals!!!

OpenStudy (anonymous):

i will help you hun

OpenStudy (anonymous):

Thanks you!!!! Choose the point-slope form of the equation below that represents the line that passes through the point (-1, 6) and has a slope of -3. a y - 6 = -3x - 3 b y - 6 = -3(x + 1) c y = -3x + 3 d 3x + y = 3

OpenStudy (anonymous):

@Tazmaniadevil

OpenStudy (anonymous):

y − 4 = −(x + 6)

OpenStudy (anonymous):

opps sorry wrong question im between to questions

OpenStudy (anonymous):

its a

OpenStudy (anonymous):

Espero que ayude u y tener un buen día hun